OpenStudy (anonymous):

Find the zeros of the function f(x) = x2 - 7x + 6 by factoring. What is the largest zero of the function? Enter a number only.

OpenStudy (dumbcow):

-6 * -1 = 6 -6 -1 = -7 --> (x-6)(x-1) = 0